Shift report 20000720 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Jul 20 2000 - 08:14:02 EDT

  • Next message: http account: "Shift report 20000720 08:00-16:00"

    ------------------------------------------------------------
    Date: Thu Jul 20 08:14:01 2000, Shift: 00:00-08:00
    Supervisor: JHL, with: ZIS
    Log book#3 pages 31 - 32
    
    Shift Summary:
    ------------------------------------------------------------
    Another quiet but not productive shift.
    
    Beam: Mainly ramping tests. Problems with 
    "booster extraction" ("experts in", "ring access",...).
    Comes and goes at the injection energy without much
    accelerations. 
    
    Data: No data is taken except pedestal runs.
    
    DAQ: "sim-comm" and "fb2-comm" red warning/errors. Couldn't
    be cleared.
    
    Detectors:  
    H2: HV Bad channels. H2T01, H2T05, H2T07 read 0. 
    Cannot hold voltages.
     
    Tile: channel#1 HV. actual reading 2V (set point: 1750V), 
    without a "red flag" in the display.  Cannot be brought up. 
    
          
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Jul 20 2000 - 08:17:20 EDT